# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v2

@@ 7 Conclusion
Our methodological approach enables us to go beyond the estimation of causal effects. Specifically, we conduct a counterfactual policy analysis to infer the consequences of adopting a series of different redistricting reforms nationwide. We find that adopting redistricting commissions generally reduces the current Republican advantage, but commissions are more effective at reducing bias when they place greater restrictions on partisan actors. For example, a Michigan-style reform (which combines a nonpartisan commission with no partisan veto points) would yield a much greater pro-Democratic effect than commissions adopted in Ohio and New York (which maintain the potential for some partisan control). While we apply this approach to three commission reforms, the same model can be applied to other reforms of interest.
@@
[…]
@@
[…]
@@
[…]
@@
$\begin{cases*}\texttt{stale\_slope}\cdot x-\texttt{out\_nonp\_part\_adv}&if Court Control is coded as ``Democrats''\\ \texttt{stale\_slope}\cdot x+\texttt{out\_nonp\_part\_adv}&if Court Control iscoded as ``Republicans''\end{cases*}$
@@
[…]
@@
[…]
@@
The first move belongs to the Republican party, which controls the state legislature. It has to pick the amount of partisan bias $-4<x<4$ in the plan that it adopts. If this plan is ultimately the final enacted plan, the Republicans receive utility $x$ and the Democrats receive utility $-x$. The second move also belongs to the Republican party, which controls the governorship. The party must decide whether or not to veto the adopted plan; if it does, there is a second round of drawing and vetoing by the legislature and governor (controlled by Republicans).
If a plan is adopted at either the first or second round, it proceeds to possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias $x^{\prime}$, then the Republicans receive utility $x^{\prime}$ and the Democrats receive utility $-x^{\prime}$. In Alabama, court review is generally not allowed on partisan grounds, but challenges under the federal Voting Rights Act are possible, so there is moderate probability of the legislature-adopted plan being overturned. If the legislature’s plan is vetoed both times, the process results in a stalemate. Since there is no enumerated stalemate procedure in Alabama, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
@@
For completeness, we consider a series of alternative outcomes which measure the partisan bias of redistricting plans. We first consider an alternative to the efficiency gap, the dilution asymmetry (Gordon and Yntiso,, 2024). This measures the difference in the percentage of each party’s votes that are wasted (cf. the number of wasted votes used in the calculation of the efficiency gap). Similar to the efficiency gap, we consider both the absolute value of differences from simulations and signed raw values for the dilution asymmetry.

### v3

@@ 7 Conclusion
Our methodological approach enables us to go beyond the estimation of causal effects. Specifically, we conduct a counterfactual policy analysis to infer the consequences of adopting a series of different redistricting reforms nationwide. We find that adopting redistricting commissions generally reduces the current Republican advantage, and this reduction is substantially larger when the reforms place greater restrictions on partisan actors. For example, a Michigan-style reform (which combines a nonpartisan commission with no partisan veto points) would yield a greater pro-Democratic effect than commissions adopted in Ohio and New York (which maintain the potential for some partisan control). While we apply this approach to three commission reforms, the same model can be applied to other reforms of interest. Future research should also apply our methodology to other redistricting problems, such as the one for state legislatures.
@@
    $\texttt{court\_outcome}(x)$ $=\frac{\texttt{out\_nonp\_bias}_{2}}{\arctan(2/2)}\cdot\arctan(x/2)+a\cdot\texttt{out\_nonp\_part\_adv}$
@@
    Notice that, in expectation, $\texttt{court\_outcome}(x)$ is increasing in $x$. This reflects our experience in following partisan gerrymandering litigation in state courts. For example, the Ohio state legislative plan was challenged multiple times in court, the outcome of each case inched towards a neutral map (according to simulations), rather than immediately jumping to a neutral map. In this case (as in many), this is because the legislature was given a new opportunity to draw a compliant plan; in practice, map-drawers tend to revise the existing plan as little as possible to stave off further litigation without abandoning the underlying partisan goals.
@@
    $\texttt{court\_vra\_outcome}(x)=\texttt{vra\_out\_slope}(x-\texttt{vra\_out\_breakeven})+\texttt{vra\_out\_breakeven};$
@@
$\texttt{exp\_court}(x)$ $=\texttt{pr\_int\_net}(x)\cdot\texttt{court\_outcome}(x)$
$\qquad+(1-\texttt{pr\_int\_net}(x))\cdot\texttt{pr\_vra\_net}(x)\cdot\texttt{court\_vra\_outcome}(x)$
$\qquad+(1-\texttt{pr\_int\_net}(x))\cdot(1-\texttt{pr\_vra\_net}(x))\cdot x$
$\texttt{pr\_int\_net}(x)$ $=\texttt{pr\_chal\_poss}\cdot\texttt{pr\_chal\_if\_poss}(x)\cdot\texttt{pr\_intervene}(x)$
$\texttt{pr\_vra\_net}(x)$ $=\texttt{pr\_vra\_chal}(x)\cdot\texttt{vra\_interv\_prob}$
@@
[…]
@@
$\texttt{pr\_veto\_chal}(x)=\texttt{veto\_nonp\_prob\_max}\cdot\texttt{pr\_chal\_if\_poss}(x),$
@@
$\texttt{exp\_drawer\_r2\_nonp}(x)$ $=x+\texttt{veto\_party}\cdot\texttt{veto\_nonp\_shift}$
@@
The first move belongs to the Republican party, which controls the state legislature. It has to pick the amount of partisan bias 4<x<4-4<x<4 in the plan that it adopts. If this plan is ultimately the final enacted plan, the Republicans receive utility $x$ and the Democrats receive utility x-x. The second move also belongs to the Republican party, which controls the governorship. The party must decide whether or not to veto the adopted plan; if it does, there is a second round of drawing and vetoing by the legislature and governor (controlled by Republicans).
If a plan is adopted at either the first or second round, it proceeds to possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias $x^{\prime}$, then the Republicans receive utility $x^{\prime}$ and the Democrats receive utility $x^{\prime}$. In Alabama, court review is generally not allowed on partisan grounds, but challenges under the federal Voting Rights Act are possible, so there is moderate probability of the legislature-adopted plan being overturned. If the legislature’s plan is vetoed both times, the process results in a stalemate. Since there is no enumerated stalemate procedure in Alabama, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
@@
This section provides additional details on the redistricting simulations used in the analyses here.
@@
The 2020 simulations were taken from McCartan et al., (2022). As that paper documents, these simulations were generated through a consistent workflow based on shared template code. Each set of sampled plans underwent manual review by two human analysts, which included checks for statistical quality based on convergence diagnostic, as well as empirical checks on aspects of the generated plans, such as demographics, county splits, compactness, etc. These are explained at greater length below.
The 2010 simulations were generated in the exact same fashion, by making modifications where appropriate to the 2020 code, and using different district apportionments and Census data. All of the code for generating the simulations for both 2010 and 2020 is available at https://github.com/alarm-redist/fifty-states. That repository also includes documentation of the specific constraints and choices made in sampling plans for each state.
[…]
  • Distribution of the final SMC weights. Highly variable or heavy-tailed weights may indicate difficulty approximating the target distribution and high variance for downstream estimates.
  • Plan diversity as measured by the variation of information metric. Identifies cases where many of the sampled plans are identical or nearly identical, which leads to discreteness artifacts and may indicate high variance and a possible lack of representativeness.
  • Gelman-Rubin $\hat{R}$ statistics for each calculated summary statistic (demographics, vote totals, compactness, county splits, etc.) These check that fully independent runs of the algorithm produce samples from the same distribution. High $\hat{R}$ values indicate that the algorithm has not converged and that the samples are not representative of their target distribution.
[…]
The values of the statistics below were also compared to those for the enacted plan. Significant discrepancies, as judged by the reviewers, could indicate a mismatch between the
  • Population deviation from the ideal population. Most states had a limit of 0.5% deviation; some, like Iowa, had a lower threshold.
  • Geographic compactness, as measured by the Polsby–Popper score or a graph-theoretic measure counting the number of removed edges. Some states used alternative measures of compactness as specified by law.
  • County and municipality splits.
  • Approximate Voting Rights Act (VRA) compliance. The number of districts with minority voting-age population fractions in certain ranges (e.g., majority-minority districts) and certain partisan ranges. If simulated maps produced fewer majority-minority districts with the same partisan characteristics as the enacted plan, then it is likely they would not have satisfied the VRA. We deferred judgement as to the required number and type of majority-minority districts to the enacting bodies.
[…]
Redistricting simulation algorithms generate samples from a particular target probability distribution. There are two main limitations of simulation approaches that stem from this fact.
First, sampling from complex, constrained spaces is a difficult statistical problem and can be done only approximately. Moreover, even if the sampling were perfect (i.e., we could generate i.i.d. samples from the target), the use of a finite set of samples to approximate the target distribution also leads to Monte Carlo errors. The potential risk of unacceptably high approximation error motivates the statistical quality control checks described above.
The second limitation is that the target distribution is designed to be both computationally feasible to sample from and to approximate the distribution of legal plans that would be drawn by a map drawer without any partisan information or bias. The default distribution that our samples are drawn from is described formally in McCartan and Imai, (2023). It is uniform on the space of all contiguous districting plans with a certain population constraint, after reweighting according to a certain graph-theoretic measure of compactness. This target distribution can be further modified, within limits, based on other features of the simulation problem. A different choice of target distribution, which interprets the legal constraints differently, or which makes a different tradeoff with computational feasibility, would lead to different samples.
Thus, there is an inherent sensitivity of the results of any simulation analysis to the choice of target distribution. This is in many ways a key advantage of simulation algorithms—the target distribution can be tuned to ask different questions about districting plans in a state. But for an analysis such as ours, it also raises questions about the sensitivity of downstream analyses to these choices.
[…]
As mentioned above, the default distribution targeted by the simulation algorithm used here is modified by the inclusion of other constraints. While some of these additional constraints can be easily described mathematically and are documented alongside the simulation code, others have effects on the target that are more difficult to express in closed form.
The first is the incorporation of county and municipality boundaries, a constraint that is applied in the vast majority of states. This changes both the support of the target (it limits the number of splits of counties or municipalities) as well as the graph-theoretic compactness measure that appears as a weighting term in the target distribution. Moreover, in many states, we used a hybrid approach where counties were protected in most of the state, but inside metropolitan areas (where many districts might appear in the same county), we used municipality boundaries instead. This allowed us to trade off county and municipality splits to better approximate the balance seen in the enacted plans. A different treatment of these administrative boundaries would lead to a different target distribution.
The second is a multi-step simulation approach used in several large states with many districts, such as Texas, Florida, and California. In these states, the simulation algorithm was first run inside a designated set of counties, such as those comprising a major metropolitan area. Rarely would the population of this area exactly divide the ideal district population, so some population in these areas was left unassigned. This first-stage simulation was repeated possibly in other areas of the state as well. Then the algorithm was run a final time to assign the remaining population to districts.
Such a multi-stage approach is supported by the design of sequential Monte Carlo algorithms, but it does introduce an implicit constraint on the target distribution, altering it from the default described in McCartan and Imai, (2023). Specifically, this simulation approach allowed for at most one district to span the border of the first-stage metropolitan areas. That is, there would be at most one district that took in the unassigned population from the first stage and connected it to population outside the first-stage area. Since the first-stage areas were defined by county borders, which are already protected by the algorithm in a similar way (i.e., only one district can span a given county border), in some states the effect of this setup may have been absent or minimal. In other states, this implicit constraint may have led to larger changes in the target distribution.
These changes and other decisions were made in an attempt to produce a sample representative of legal plans that would be drawn by a map drawer without any partisan information or bias. How might they affect the analyses conducted in this paper? A key point is that only the mean and standard deviation of the simulated distribution (rather than the entire distribution) were used to define the outcome variables in the analyses here (except the partisan harm outcomes used in the appendices only). In our experience, and as supported by the supplementary analysis conducted in Appendix J of Kenny et al., (2023), the first two moments are less sensitive to alternative simulation choices than other summary statistics, such as p-values from the tails of the distribution.
Even more importantly, the moments of the simulation distributions were not themselves the estimand of interest here. Rather, they were used to define an outcome variable, which was then time-differenced and fed into a regression model. So any biases or sensitivity in the simulation distribution that were constant over time would not affect the results. Moreover, any such sensitivity that is uncorrelated with the explanatory variables, after differencing, would likewise have no effect. If sensitivity to simulation setup is considered as a form of measurement error, then the fact that simulations are used only in the regression outcome means that there is little risk of attenuation bias as there might be for predictor variables that use simulations.
[…]
For completeness, we consider a series of alternative outcomes which measure the partisan bias of redistricting plans. We start with the efficiency gap, calculated as the difference between the wasted votes of the two parties (Stephanopoulos and McGhee,, 2015). Wasted votes are votes cast for losing candidates and votes cast for winning candidates beyond the threshold needed to win. The idea behind the efficiency gap is that a fair plan should have similar numbers of wasted votes for both parties. The partisan-signed version of the efficiency gap indicates which party benefits from the plan in terms of vote efficiency; we use positive values to indicate pro-Republican bias. Unlike with Republican seat share, we do not subtract the average simulated value from the partisan-signed efficiency gap measure. This is because the efficiency gap is already designed to have a universal scale, where 0 indicates partisan fairness; subtracting the simulated average would change the definition of a fair baseline for this measure. To create a nonpartisan measure, as with Republican seats, we take the absolute value of the difference between the enacted and average simulated plan’s efficiency gap. Smaller values indicate a fairer plan by this measure (closer to the normatively preferred value of 0).
Despite its popularity in the redistricting literature and litigation, the efficiency gap faces a number of important limitations. For example, the definition of the efficiency gap implies an ideal allocation of evenly wasted votes between parties, which can be achieved through a series of 75-to-25 districts. This “three-to-one” promotion has been challenged as an unreasonable standard for fairness (Bernstein and Walch,, 2022). Further, because the efficiency gap calculation is based in part on the number of seats, states with small numbers of seats can have near discrete, “lumpy” distributions of the efficiency gap (Nagle,, 2017; Cho,, 2017). Finally, this seat-based measure means that relatively small changes in vote share can cause large changes in the efficiency gap, when those changes in vote share cause seat control to flip (Stephanopoulos and McGhee,, 2018).
@@
We account for these limitations by considering a series of alternative measures, including the dilution asymmetry (Gordon and Yntiso,, 2024). This measures the difference in the percentage of each party’s votes that are wasted (cf. the number of wasted votes used in the calculation of the efficiency gap). Similar to the efficiency gap, we consider both the absolute value of differences from simulations and signed raw values for the dilution asymmetry.
@@
[…]
The setup of the linear model used for the primary effect estimates is described in the estimation section in the main text. In terms of the prior specification, the intercept receives a weakly informative prior of $\mathcal{N}\quantity(0,(2.5\cdot\sigma_{y})^{2})$, where $\sigma_{y}^{2}$ is the variance of $\Delta Y_{t}$. Each main effect coefficient receives a weakly informative prior of $\mathcal{N}\quantity(0,(0.75\cdot\sigma_{y}/\sigma_{x})^{2})$, where $\sigma_{x}^{2}$ is the variance of each covariate $x$. The prior standard deviation is further scaled by 0.25 for interaction terms, following the discussion in Gelman et al., (2008). Finally, the residual standard deviation receives an $\operatorname{Expo}(1/\sigma_{y})$ prior. Computation is carried out via NUTS-HMC as implemented in the Stan probabilistic programming language (Stan Development Team,, 2024).
@@
[…]
This paper focuses on the plans used in the first election after redistricting, whereas litigation can change the enacted plan well after this point. As such, plans which are labelled as having been drawn by legislatures and, less frequently, commissions, may eventually be redrawn by courts. Peeking at such eventualities in designing our game theoretic model can introduce post-treatment bias, so we avoid this. Notably, states in the South are susceptible to VRA litigation that can take more than the 14 months between Census data release and November general elections. They may see more restriction in the model than is empirically true in that time period.
Below, we demonstrate that tuning parameters for solving the game do not meaningfully impact the calculated Nash equilibria. To produce this, we lower the model parameters across the board. We largely decrease the likelihood of VRA litigation changing the final plan and modestly decrease the probability that other litigation changes the final plan.
[…]